OpenStudy (anonymous):

Find the inverse of the function. f(x) = x^3 - 3

OpenStudy (psymon):

If you have y = x^3 - 3, then you need to switch y and x with each other, so you end up with x = y^3 - 3. Then just solve for y.

OpenStudy (anonymous):

f(x)=y= x^3 - 3 swap x and y and solve for y :)

OpenStudy (anonymous):

thanks both of you (: